# OPTiML: Dense Semantic Invariance Using Optimal Transport for Self-Supervised Medical Image Representation

### v1

@@ Abstract
Self-supervised learning (SSL) has emerged as a promising technique for medical image analysis due to its ability to learn without annotations. However, despite the promising potential, conventional SSL methods encounter limitations, including challenges in achieving semantic alignment and capturing subtle details. This leads to suboptimal representations, which fail to accurately capture the underlying anatomical structures and pathological details. In response to these constraints, we introduce a novel SSL framework OPTiML, employing optimal transport (OT), to capture the dense semantic invariance and fine-grained details, thereby enhancing the overall effectiveness of SSL in medical image representation learning. The core idea is to integrate OT with a cross-viewpoint semantics infusion module (CV-SIM), which effectively captures complex fine-grained details inherent in medical images across different viewpoints. In addition to the CV-SIM module, OPTiML imposes the variance and covariance regularizations within OT framework to force the model focus on clinically relevant information while discarding less informative features. Through these, the proposed framework demonstrates its capacity to learn semantically rich representations that can be applied to various medical imaging tasks. To validate its effectiveness, we conduct experimental studies on three publically available datasets from chest X-ray modalitiy. Our empirical results reveal OPTiML’s superiority over state-of-the-art methods, across all evaluated tasks.
Index Terms:
Self-supervised learning, Medical image analysis, Chest X-ray, Representation learning, Optimal Transport
I Introduction
@@
Self-supervised learning (SSL) has gained significant momentum in the field of deep learning over the recent years [mocov2, simclr, barlow, byol, gorade2023large]. In the medical image analysis domain, where large-scale annotated datasets are challenging to obtain, SSL is particularly useful due to its ability to harness unlabeled data [medaug, mococxr, micle]. However, conventional SSL methods, like SimCLR [simclr], MoCo [mococxr], BYOL [byol], Barlow-Twins [barlow], etc., though effective in various domains, encounter challenges when applied to the subtle complexities of medical images. The interpretation of medical images necessitates the deep understanding of both anatomical structures and pathological indicators for precise diagnosis and treatment decisions. This requirement poses a significant challenge for SSL methods that focus on general feature learning, while, often ignoring the fine-grained details and clinically relevant patterns [multi-gcn, synergynet, semiGCN, micle, pacl].
@@
Conventional SSL methods treat the entire image as a collection of isolated features, thereby missing out on the context and spatial dependencies between different anatomical regions. Moreover, the fine-grained variations in medical images are particularly useful due to the coexistence of anatomical heterogeneity and the visual similarity of different regions. Heterogeneity pertains to the visual distinctions between instances of the same disease or anatomical structures [zheng2021ressl]. For instance, the shape and size of a nodule may vary based on its stage and location. Conversely, visual similarity refers to the challenges in distinguishing between different diseases or structures that share common visual characteristics [xing2021categorical]. For instance, consolidation and pneumonia are labeled separately even though they are the types of lung opacities and have a natural relation. In context of SSL for medical applications it becomes imperative for models to adeptly capture these subtle variations. Failure to do so may result in trivial solutions and hinder the acquisition of a representative feature space, essential for meaningful medical image analysis [zhao2022cross, wang2022rcl].
@@
In addressing these challenges, the concept of dense semantic invariance holds potential to improve SSL for medical image analysis. It empowers the model to capture precise anatomical and pathological attributes, regardless of differences in viewpoints, orientations, and imaging conditions. In this context, we propose OPTiML, a novel solution that integrates concepts from SSL and optimal transport (OT) [luise2018differential] to foster dense semantic invariance in medical image representation learning. The application of OT, provides a robust mechanism to measure the dissimilarity between representations and to find an optimal mapping between them. Additionally, OPTiML incorporates the cross-viewpoint semantics infusion module (CV-SIM) to enhance the model’s capability in capturing fine details from diverse viewpoints, thereby improving the alignment of semantically relevant features. Furthermore, to ensure that the acquired representations are both informative and less redundant, we apply variance and covariance regularizations [vicreg] during the representation learning process. By incorporating variance regularization, OPTiML ensures stability in the learned representations, preventing the model from overemphasizing noise and irrelevant features. Concurrently, covariance regularization helps in preserving relationships and dependencies between different features within the images. This is particularly beneficial for capturing the intricate spatial and contextual relationships among anatomical structures or pathological patterns. The integration of variance and covariance regularizations within the OT framework enhances the ability of model to generate stable and semantically aligned representations.
The overall contributions of this paper are summarized as follows:
  • OPTiML introduces an integration of OT into the SSL framework to achieve dense semantic invariance in medical image representations.
  • OPTiML incorporates a novel CV-SIM module, which enhances the model’s ability to capture fine details from diverse viewpoints, thereby contributing to a more comprehensive and accurate representations.
  • We evaluate the performance of OPTiML on diverse chest X-ray datasets, including NIH-Chest X-ray14, Vinbig-CXR, and RSNA. The results highlight effectiveness of the proposed model with consistent performance improvements across various medical imaging tasks.
In the subsequent sections of this paper, we provide a detailed exposition of the components of the OPTiML framework, elucidate the theoretical underpinnings of the CV-SIM module, delve into the OT application in the context of medical image representation, and present the comprehensive experimental analysis that validates the efficacy of our approach.
@@
Figure 1: Architecture of the OPTiML Framework. $I_{s}$ and $I_{t}$ are two augmented versions of $I$, pass through backbone encoders $f_{s}$ and $f_{t}$ to obtain $z_{s}$ and $z_{t}$ which OT solver subsequently utilizes along with the cost matrix $M$. $g_{s}$ and $g_{t}$ are the feature vectors after the global average pooling (GAP) layer which are expanded to $q_{s}$ and $q_{t}$, to increase representational variability.
II Background and Related Works
@@
Self-supervised Learning. The current SSL approaches employ a joint embedding architecture to acquire representations that are invariant to diverse perspectives [simclr, byol, vicreg]. These techniques have varying ways of circumventing the collapse of solutions. Negative sampling is employed by contrastive methods [simclr, mocov2, mococxr, medaug] to push dissimilar samples apart from each other while clustering methods such as proposed in [SwaV] ensure an equitable distribution of samples within clusters. Non-contrastive techniques [byol, simclr, barlow, vicreg, dino], which serve as the alternative of contrastive methods, maintain the representations’ information content either by adopting architectural constraints such as asymmetric network, momentum encoder, stop-gradient, etc. or by applying specific regularization [barlow, vicreg]. Unlike the aforementioned global methods, local methods concentrate on learning a group of local features that describe smaller segments of an image. For instance, in [pixelpro] authors addressed the challenge of pixel-level self-supervised representation learning, crucial for tasks like object detection and semantic segmentation by leveraging pixel-level contrastive learning. Similarly DenseCL [denseCL], a novel dense contrastive learning paradigm, introduced pixel-level SSL using a dense projection head and dense contrastive loss.
@@
SSL for Medical Image Analysis. SSL has become an increasingly popular approach for medical image analysis [shurrab2022self]. A majority of recent works have shown the effectiveness of SSL for tasks such as segmentation, classification, and localization [zhang2022dive, mococxr, medaug, huang2021lesion].The adoption of SSL techniques has predominantly leaned towards the utilization of contrastive learning approaches. However, in stark contrast, the number of approaches embracing non-contrastive SSL within this domain remains notably limited. For instance in [cho2023chess], authors introduced CheSS, a publicly accessible model pre-trained with 4.8-M chest X-ray images through contrastive learning on a large dataset. In [micle], the authors presented a distinct SSL approach termed Multi-Instance Contrastive Learning (MICLe), which harnesses the multiple studies of the subject to create positive pairs. In another study [dira], authors proposed a collaborative SSL framework DiRA that unites discriminative, restorative, and adversarial learning to learn fine-grained representations. MedAug [medaug] and PaCL [pacl] are another works based upon the contrastive approach that incorporates insightful metadata for selective positive and negative pairs. PCRL [pcrlv1, pcrlv2], another contrastive approach, that enhances medical image representations by dynamically reconstructing diverse image contexts. This approach is applicable to both X-ray and CT modalities.
@@
Optimal Transport Theory. OT determines the most efficient way to transform one distribution into another, considering the associated costs or distances between individual elements in the distributions [chen2021wasserstein]. It is popular for computer vision tasks like domain adaptation [ding2023cross], multitask learning [janati2019wasserstein], feature matching [sarlin2020superglue], generative model [bunne2019learning] etc. where measuring dissimilarity or similarity between distributions is essential. Zheng et al.  [ge2021ota] framed label assignment in object detection, as an OT problem where the unit transportation cost between anchor and ground-truth pairs was defined by weighted summation of classification and regression losses, solved through Sinkhorn-Knopp Iteration for optimal assignment. Kim et al. proposed an identity-invariant facial expression recognition method using OT to quantify inter-identity variation, for optimal matching of similar expressions across different identities [kim2022optimal]. Zhu et al. introduced OT-guided translation network for unpaired image-to-image translation scheme to enhance retinal color fundus photography [zhu2023otre]. CUI et .al in [yufei2023retrieval] leveraged OT for nearest neighbor retrieval in on whole slide images (WSIs) to address performance degradation in multiple instance learning models when tested on out-of-domain data. OT has limited applicability in the medical domain due to complexities inherent in medical image data structures. In this work, we introduce a novel integration of OT within the SSL framework to enhance the effectively handle the medical image tasks.
III Proposed Method
@@
In this section we first present preliminaries on OT theory, subsequently describe the proposed SSL framework OPTiML which comprises of two main stages: 1) formulation of dense semantic invariance as an OT problem and 2) cross-viewpoint semantics infusion module (CV-SIM).
@@
OPTiML SSL Framework. OT has recently gained interest in computer vision applications which involve comparison of the images. By determining the optimal transportation plan between two images, we can quantify the similarity or dissimilarity between them in a way that considers the spatial arrangement of pixel intensities. This is particularly useful in applications like medical images where the pixel-wise differences may not capture perceptual similarity.
@@
In this work we adapt OT in SSL to measure the optimal distance between two sets of image features which represent the embeddings extracted from each image. The primary objective is to ascertain the optimal flow of information between source and target images while ensuring the minimum associated cost in terms of distance between features. More specifically, as illustrated in Figure 1 OPTiML framework generates two augmented versions ($I_{s}$ and $I_{t}$) of the original input image $I$ by applying random augmentations $t_{s}$ and $t_{t}$ sampled from a predefined set of augmentations $\mathcal{T}$. $I_{s}$ and $I_{t}$ undergo processing through two weight-shared CNN encoders $f_{s}$ and $f_{t}$. The output of $f_{s}$ and $f_{t}$ are the dense feature maps $z_{s}\in\mathbb{R}^{d\times hw}$ and $z_{t}\in\mathbb{R}^{d\times hw}$ respectively, where $d$ represents the number of channels while $hw$ captures the spatial dimensions. Subsequently $z_{s}$ and $z_{t}$ are utilized to compute the discrepancy matrix $C_{ij}\in\mathbb{R}^{d\times d}$, where each element quantifies the cosine similarity between feature vectors at position $i$ in $z_{s}$ and the feature vector at position $j$ in $z_{t}$. The discrepancy matrix $C_{ij}$ is constructed as follows:
@@
Dense feature maps $z_{s}$ and $z_{t}$ offer advantage over post-pooling feature vectors for semantic alignment, as they preserve the spatial information and relationships between pixels in the original image. This enables the model to learn more meaningful representations for medical image applications, where the spatial arrangement of structures is vital for accurate diagnosis. The formulation of the OT problem aligns seamlessly with dense feature maps $z_{s}$ and $z_{t}$, allowing for a more granular and accurate estimation of the cost associated with transferring of information between corresponding positions.
@@
Furthermore, $z_{s}$ and $z_{t}$ undergo a global average pooling layer $GAP$ to output $g_{s}$ and $g_{t}$, that condenses the spatial information captured by the dense feature maps, into single-dimensional compact representations. Subsequently, $g_{s}$ and $g_{t}$, are projected to an high dimensional space using $MLP$ expansion heads $h_{s}$ and $h_{t}$ to output $q_{s}$ and $q_{t}$ respectively. $g_{s}$ and $g_{t}$ are further utilized by the CV-SIM modules $R_{s}$ and $R_{t}$ in each branch of the network. The subsequent section provides the details of the CV-SIM.
@@
Cross-Viewpoint Semantics Infusion Module. CV-SIM module incorporates, multi-head cross-view attention mechanism, where each head attends to diverse subspaces of queries and keys. This contributes in navigating the high-dimensional feature space to extract subtle relationships and dependencies across different viewpoints within the data. In this context, it linearly transform $g_{s}$ into $q$ and $g_{t}$ into $k$ and $v$, each having a dimension of $\mathbb{R}^{d}$. Subsequently, a multi-head attention mechanism computes attention scores $\hat{a}$ through a scaled dot-product between queries ($q$) and keys ($k$) by employing $\sigma(\frac{q.k^{T}}{\sqrt{dim}})$. $\sigma$ represents the softmax function, which transforms relevance attention scores into attention weights while $dim$ denotes the dimensionality of the query and key vector.
@@
Subsequently, CV-SIM, further refines the representation by computing the attended values, through a weighted sum operation, where the attention weights guide the aggregation of information from the target representation $v$ based on the computed attention weights $\hat{a}$. This, ensures that CV-SIM captures intricate dependencies across different viewpoints, contributing to the enrichment of the overall feature space. The attended values computation yields an enhanced representation $r_{s}\in\mathbb{R}^{d}$. Likewise, a corresponding procedure is executed in $R_{t}$, where the roles of $g_{s}$ and $g_{t}$ are interchanged to generate the corresponding outcome $r_{t}\in\mathbb{R}^{d}$. The resulting enhanced representations $r_{s}$ and $r_{t}$ from $R_{s}$ and $R_{t}$ respectively, undergoes another transformation through the softmax function to convert it into a probabilistic distribution $\mu$ and $\nu$ over the visual feature spaces respectively. OT solver utilized these $\mu$ and $\nu$ as the marginal constraints of the transport plan such that the total cost of transporting features from one image to another is minimum. This step ensures that the probability distribution derived from the refined representation is effectively incorporated into the optimization process facilitated by the OT solver, contributing to the alignment of features across different branches within the OPTiML framework.
TABLE I: Comparison of the considered approaches under the semi-supervised linear evaluation protocol for the classification task on NIH, Vinbig-CXR and RSNA datasets, after finetuning the representations using different subsets of labeled samples from training set. The arrows $\uparrow$ and $\downarrow$ indicate improvement and decrement compared with bs.

### v4

# OTCXR: Rethinking Self-supervised Alignment using Optimal Transport for Chest X-ray Analysis

@@ Abstract
Self-supervised learning (SSL) has emerged as a promising technique for analyzing medical modalities such as X-rays due to its ability to learn without annotations. However, conventional SSL methods face challenges in achieving semantic alignment and capturing subtle details, which limits their ability to accurately represent the underlying anatomical structures and pathological features. To address these limitations, we propose OTCXR, a novel SSL framework that leverages optimal transport (OT) to learn dense semantic invariance. By integrating OT with our innovative Cross-Viewpoint Semantics Infusion Module (CV-SIM), OTCXR enhances the model’s ability to capture not only local spatial features but also global contextual dependencies across different viewpoints. This approach enriches the effectiveness of SSL in the context of chest radiographs. Furthermore, OTCXR incorporates variance and covariance regularizations within the OT framework to prioritize clinically relevant information while suppressing less informative features. This ensures that the learned representations are comprehensive and discriminative, particularly beneficial for tasks such as thoracic disease diagnosis. We validate OTCXR’s efficacy through comprehensive experiments on three publicly available chest X-ray datasets. Our empirical results demonstrate the superiority of OTCXR over state-of-the-art methods across all evaluated tasks, confirming its capability to learn semantically rich representations.
1 Introduction
@@
Self-supervised learning (SSL) has gained significant momentum in deep learning in recent years [mocov2, simclr, barlow, byol, gorade2023large]. In the medical image analysis domain, where large-scale annotated datasets are challenging to obtain, SSL is particularly useful due to its ability to harness unlabeled data [medaug, mococxr, micle]. However, conventional SSL methods, such as SimCLR [simclr], MoCo [mococxr], BYOL [byol], Barlow-Twins [barlow], etc., though effective in various domains, encounter challenges when applying to chest X-rays(CXRs). The interpretation of CXRs necessitates a deep understanding of anatomical structures and pathological indicators for precise diagnosis and treatment decisions. This requirement poses a significant challenge for SSL methods that focus on generic feature learning while often discarding the clinically relevant patterns [multi-gcn, synergynet, semiGCN, micle, pacl].
@@
Conventional SSL methods treat the entire image as a collection of isolated features, thereby missing local information and global contextual dependencies between different anatomical regions. Moreover, local variations in CXRs are particularly useful because of the coexistence of anatomical heterogeneity and visual similarity of different regions. Heterogeneity refers to visual distinctions between instances of the same disease or anatomical structures [zheng2021ressl]. For instance, the shape and size of a nodule may vary based on its stage and location. Conversely, visual similarity refers to the challenges in distinguishing between different diseases or structures that share common visual characteristics [xing2021categorical]. For instance, consolidation and pneumonia are labeled separately even though they are the types of lung opacities and have a natural relation. In the context of SSL pre-training for CXRs, it becomes essential for models to capture these local variations. Failure to do so may result in trivial solutions and hinder the acquisition of representative feature space, essential for meaningful CXR analysis [zhao2022cross, wang2022rcl].
@@
Addressing these challenges, we propose rethinking the concept of alignment through invariance by introducing the concept of dense semantic invariance, potentially improving SSL pre-training in the context of CXRs. It empowers the model to capture precise anatomical and pathological attributes regardless of differences in viewpoints, orientations, and imaging conditions. To this end, we propose OTCXR, a novel objective that replaces existing invariance-based alignment strategies in SSL with dense semantic invariance, modeled using Optimal Transport Theory (OT) [luise2018differential]. Further, the OTCXR integrates the Cross-Viewpoint Semantics Infusion Module (CV-SIM) to enhance the model’s ability to capture global contextual dependencies across diverse viewpoints. The OT solver utilizes the resulting CV-SIM representations as marginal constraints in the transport plan, minimizing the total cost of transporting features between images. This approach effectively improves the alignment of semantically relevant features. Furthermore, to ensure that the acquired representations are informative and less redundant, we apply variance and covariance regularizations [vicreg] during the representation learning process. The integration of variance and covariance regularization within the OT framework enhances the ability of the model to generate stable and semantically aligned representations. Our contributions are summarized as follows:
  • We introduce OTCXR, a novel SSL framework that formulates dense semantic invariance as an OT problem, enabling fine-grained feature alignment effective for chest X-ray analysis during pre-training.
  • We propose a Cross-Viewpoint Semantics Infusion Module (CV-SIM) that complements the OT approach by capturing global semantic relationships, enhancing the framework’s ability to learn clinically relevant features across different views and transformations.
  • Through comprehensive evaluations on diverse chest X-ray datasets (NIH-Chest X-ray14, Vinbig-CXR, and RSNA), we demonstrate OTCXR’s effectiveness in various medical imaging tasks, particularly in limited data scenarios, showcasing consistent performance improvements over existing SSL methods.
@@
Figure 1: Architecture of the OTCXR Framework. $I_{s}$ and $I_{t}$ are two augmented versions of $I$, pass through backbone encoders $f_{s}$ and $f_{t}$ to obtain $z_{s}$ and $z_{t}$ which OT solver subsequently utilizes along with the cost matrix $M$. $g_{s}$ and $g_{t}$ are the feature vectors after the global average pooling (GAP) layer, which are expanded to $q_{s}$ and $q_{t}$ to increase representational variability.
2 Background and Related Works
@@
Self-supervised Learning. The current SSL approaches employ a joint embedding architecture to acquire representations invariant to diverse perspectives [simclr, byol, vicreg]. These techniques have varying ways of circumventing the collapse of solutions. Negative sampling is employed by contrastive methods [simclr, mocov2, mococxr, medaug] to push dissimilar samples apart from each other while clustering methods such as proposed in [SwaV] ensure an equitable distribution of samples within clusters. Non-contrastive techniques [byol, simclr, barlow, vicreg, dino], which serve as the alternative of contrastive methods, maintain the representations’ information content either by adopting architectural constraints such as asymmetric network, momentum encoder, stop-gradient, etc. or by applying specific regularization [barlow, vicreg]. Unlike the aforementioned global methods, local methods concentrate on learning a group of local features that describe smaller segments of an image. For instance, the authors in  [pixelpro] addressed the challenge of pixel-level self-supervised representation learning, crucial for tasks like object detection and semantic segmentation by leveraging pixel-level contrastive learning. Similarly, DenseCL [denseCL], a novel dense contrastive learning paradigm, introduced pixel-level SSL using a dense projection head and dense contrastive loss.
@@
SSL for CXR Analysis. SSL has become an increasingly popular approach for CXRs [shurrab2022self]. Most recent works have shown the effectiveness of SSL for tasks such as segmentation, classification, and localization [zhang2022dive, mococxr, medaug, huang2021lesion, 10666966]. The adoption of SSL techniques has predominantly leaned towards utilizing contrastive learning approaches. However, in stark contrast, the number of approaches embracing non-contrastive SSL within this domain remains notably limited. For instance, the authors in [cho2023chess] introduced CheSS, a publicly accessible model pre-trained with 4.8-M chest X-ray images through contrastive learning on a large dataset. Multi-Instance Contrastive Learning (MICLe), is another approach presented in [micle], harnesses the multiple studies of the subject to create positive pairs. In another study, authors proposed a collaborative SSL framework, DiRA, that unites discriminative, restorative, and adversarial learning to learn fine-grained representations [dira]. MedAug [medaug] and PaCL [pacl] are other works based upon the contrastive approach that incorporates insightful metadata for selective positive and negative pairs. PCRL [pcrlv1, pcrlv2], another contrastive approach that enhances representations by dynamically reconstructing diverse image contexts. This approach applies to both X-ray and CT modalities.
@@
Optimal Transport Theory. OT determines an efficient way to transform one distribution into another, considering the associated costs or distances between individual elements in the distributions [chen2021wasserstein]. It is popular for computer vision tasks like domain adaptation [ding2023cross], multitask learning [janati2019wasserstein], feature matching [sarlin2020superglue], generative model [bunne2019learning] etc. where measuring dissimilarity or similarity between distributions is essential. Zheng et al.  [ge2021ota] framed label assignment in object detection as an OT problem where the unit transportation cost between anchor and ground-truth pairs was defined by weighted summation of classification and regression losses, solved through Sinkhorn-Knopp iteration for optimal assignment. Kim et al. proposed an identity-invariant facial expression recognition method using OT to quantify inter-identity variation for optimal matching of similar expressions across different identities [kim2022optimal]. Zhu et al. introduced an OT-guided translation network for unpaired image-to-image translation schemes to enhance retinal color fundus photography [zhu2023otre]. CUI et .al in [yufei2023retrieval] leveraged OT for nearest neighbor retrieval in on whole slide images (WSIs) to address performance degradation in multiple instance learning models when tested on out-of-domain data. OT has limited applicability in the CXRs and medical domain due to complexities inherent in medical image data structures. In this work, we introduce a novel integration of OT within the SSL framework to enhance the effective handling of the CXR tasks.
3 Proposed Method
@@
In this section, we first present preliminaries on OT theory and subsequently describe the proposed SSL framework OTCXR, which comprises two main stages: 1) formulation of dense semantic invariance as an OT problem and 2) cross-viewpoint semantics infusion module (CV-SIM).
@@
OTCXR SSL Framework. OT has recently gained interest in computer vision applications that involve comparing images. By determining the optimal transportation plan between two images, we can quantify the similarity or dissimilarity between them in a way that considers the spatial arrangement of pixel intensities. This is particularly useful in CXR analysis, where the pixel-wise differences may not capture perceptual similarity.
@@
In this work, we adapt OT in SSL to measure the optimal distance between two sets of features representing the embeddings extracted from each image. The primary objective is to ascertain the optimal flow of information between source and target images while ensuring the minimum associated cost in terms of distance between features. More specifically, as illustrated in Figure 1, the OTCXR framework generates two augmented versions ($I_{s}$ and $I_{t}$) of the original input image $I$ by applying random augmentations $t_{s}$ and $t_{t}$ sampled from a predefined set of augmentations $\mathcal{T}$. Following the literature [simclr, mococxr, byol], we adopted data augmentations common in SSL pre-training. Particularly OTCXR employs random resized cropping, color jitter, random horizontal flipping, and Gaussian blur.
$I_{s}$ and $I_{t}$ undergo processing through two weight-shared CNN encoders $f_{s}$ and $f_{t}$. The output of $f_{s}$ and $f_{t}$ are the dense feature maps $z_{s}\in\mathbb{R}^{d\times hw}$ and $z_{t}\in\mathbb{R}^{d\times hw}$ respectively, where $d$ represents the number of channels while $hw$ captures the spatial dimensions. Subsequently, $z_{s}$ and $z_{t}$ are utilized to compute the discrepancy matrix $C_{ij}\in\mathbb{R}^{d\times d}$, where each element quantifies the cosine similarity between feature vectors at position $i$ in $z_{s}$ and the feature vector at position $j$ in $z_{t}$. The discrepancy matrix $C_{ij}$ is constructed as follows:
@@
Dense feature maps $z_{s}$ and $z_{t}$ offer an advantage over post-pooling feature vectors for semantic alignment, as they preserve the spatial information and relationships between pixels in the original image. This enables the model to learn more meaningful representations for CXRs, where the spatial arrangement of structures is vital for accurate diagnosis. The formulation of the OT problem aligns seamlessly with dense feature maps $z_{s}$ and $z_{t}$, allowing for a more granular and accurate estimation of the cost associated with transferring information between corresponding positions.
@@
Furthermore, $z_{s}$ and $z_{t}$ undergo a global average pooling layer $GAP$ to output $g_{s}$ and $g_{t}$, which condenses the spatial information captured by the dense feature maps into single-dimensional compact representations. Subsequently, $g_{s}$ and $g_{t}$ are projected to a high dimensional space using $MLP$ expansion heads $h_{s}$ and $h_{t}$ to output $q_{s}$ and $q_{t}$ respectively. $g_{s}$ and $g_{t}$ are further utilized by the CV-SIM modules $R_{s}$ and $R_{t}$ in each branch of the network. The subsequent section provides the details of the CV-SIM.
@@
Cross-Viewpoint Semantics Infusion Module. CV-SIM module incorporates a multi-head cross-view attention mechanism, where each head attends to diverse subspaces of queries and keys. This contributes to navigating the high-dimensional feature space to extract subtle relationships and dependencies across different viewpoints within the data. In this context, it linearly transform $g_{s}$ into $q$ and $g_{t}$ into $k$ and $v$, each having a dimension of $\mathbb{R}^{d}$. Subsequently, a multi-head attention mechanism computes attention scores $\hat{a}$ through a scaled dot-product between queries ($q$) and keys ($k$) by employing $\sigma(\frac{q.k^{T}}{\sqrt{dim}})$. $\sigma$ represents the softmax function, which transforms relevant attention scores into attention weights, while $dim$ denotes the dimensionality of the query and key vector.
@@
Subsequently, CV-SIM refines the representation by computing the attended values through a weighted sum operation, where the attention weights guide the aggregation of information from the target representation $v$ based on the computed attention weights $\hat{a}$. This ensures that CV-SIM captures intricate dependencies across different viewpoints, contributing to the enrichment of the overall feature space. The attended values computation yields an enhanced representation $r_{s}\in\mathbb{R}^{d}$. Likewise, a corresponding procedure is executed in $R_{t}$, where the roles of $g_{s}$ and $g_{t}$ are interchanged to generate the corresponding outcome $r_{t}\in\mathbb{R}^{d}$. The resulting enhanced representations $r_{s}$ and $r_{t}$ from $R_{s}$ and $R_{t}$, respectively, undergo another transformation through the softmax function to convert it into a probabilistic distribution $\mu$ and $\nu$ over the visual feature spaces respectively.
OT solver utilized this $\mu$ and $\nu$ as the marginal constraints of the transport plan such that the total cost of transporting features from one image to another is minimal. This step ensures that the probability distribution derived from the refined representation is effectively incorporated into the optimization process facilitated by the OT solver, contributing to the alignment of features across different branches within the OTCXR framework.
Table 1: Performance comparison of SSL methods and OTCXR for chest X-ray classification on NIH, VinBig-CXR, and RSNA datasets. Results show AUC scores (%) for NIH and VinBig-CXR and accuracy (%) for RSNA under semi-supervised linear evaluation, using 1%, 10%, and 100% of labeled training data. Values in parentheses indicate performance differences from the supervised baseline (Sup.2). Arrows $\uparrow$ and $\downarrow$ denote improvement or decrease, respectively; ‘-’ indicates no change.